OpenStudy (anonymous):

Bonnie has two similar rectangular boxes. The dimensions of box 1 are twice those of box 2. How many times greater is the volume of box 1 than the volume of box 2?

OpenStudy (anonymous):

Va=a1*a2*a3 Vb=b1*b2*b2 but each a is 2ce as big as b

OpenStudy (anonymous):

So: Va=2(b1)*2(b2)*2(b3)=8*b1*b2*b3=8*Vb

OpenStudy (anonymous):

Thus, the volume is 8 times greater.

OpenStudy (anonymous):

thanks Brutal

OpenStudy (anonymous):

L * W * H = 2L * 2 W * 2H = 2^3 ( L * W * H) => V1 = 8 V2